What mass (kg) of the solvent water should be added to 429 g of the solute tetrahydrofuran to generate a solution that contains 8.70 % by mass C4H8O.

Molar Mass (g/mol)
- C4H8O 72.12
- H2O 18.015
Density (g/mL)
- C4H8O 0.8892
- H2O 0.9982
Name/Formula
- tetrahydrofuran / C4H8O
- water / H2O
Answer #1
Let mass of solvent= m g ; Mass of solution= m+429 g

429/(m+429)= 8.7/100 =.087

.087m+ 429x.087 =429 ; m= 4502.03 g = 4.502 kg
Answer #2
Step1 Let mass of solvent= m g ; Mass of solution= m+429 g

Step2 429/(m+429)= 8.7/100 =.087

Step3 .087m+ 429x.087 =429 ; m= 4502.03 g = 4.502 kg
Answer #3

=> percent by mass of the solute = {m(solute)/[m(solute) + m(solvent)] }* 100

8.7 = 429/(429+m) * 100

=> m = 42900/8.7 - 429

= 4502.03 gms